Sol Summary – December 4th

Sol: 7
Summary Title: Full Power to the Suit Fans
Author’s name: Matt Eby
Mission Status: In-Simulation
Sol Activity Summary: Resolve EVA suit battery anomaly, excursion in local area to complete drone surveys and to deploy the HAM radio antenna.
Look Ahead Plan: Soil sample from Southern quadrant, second interim mirror inspection.
Anomalies in work: None
Weather: Calm winds throughout. Largely clear skies with moderate temperatures.
Crew Physical Status: Crew is in good shape, nothing of concern to note.
EVA: EVA led by the Technology Officer, with Commander and GreenHab Officer, traveled around the HAB and out to the mirror experiment location. 1 1/2 hour duration on foot.
Reports to be filed: Journalist, Commander, Operations, Photos, EVA report, EVA request, GreenHab
Support Requested: N/A

EVA Report – December 4th

EVA #9

Author: Kristine Ferrone (CDR) / Trevor Jahn (Technology Officer) / Matthew Eby (Green Hab Officer)

Purpose of EVA: Deploy ham radio antenna near Hab, visual inspection of mirror coating experiment samples, drone flights around Hab and mirror coating experiment sample location across Cow Dung Rd from Marble Ritual
Start Time: 12:00 PM MT
End Time: 1:36 PM MT

Narrative: CDR (Kristine Ferrone), Technology Officer (Trevor Jahn), and Green Hab Officer (Matthew Eby) left the airlock at 12:00 PM MT. Drone Flight Plan 1 had a successful takeoff Southeast of the Crew Habitat and conducted a successful fly over pattern over the Crew Habitat, while taking timelapse photos for later processing using photogrammetry. The crew then proceeded to the north side of the Hab and drove in the ground spike for the ham radio antenna. The base was threaded onto the ground spike, the ground wire laid out, and the extension pole and whip antenna raised for ham radio comms. The coax cable was laid out from the antenna to the side of the tunnel outside the rear airlock and IVA crew retrieved the cable from the tunnel and ran it to the ham radio station. The crew then traversed on foot to the location of the mirror coating experiment samples across Cow Dung Rd. from Marble Ritual and visually inspected the payload. Drone Flight Plan 2 took off successfully near the Mirror Sample location, and conducted 90% of its planned flight pattern, before returning to its takeoff point due to the battery falling below 25%. Timelapse photos were captured for later processing using photogrammetry, and finally, the crew traversed on foot back to the Hab.

Coordinates (use UTM WGS 84): 4251000, 518500

Participants: CDR (Kristine Ferrone), Technology Officer (Trevor Jahn), Green Hab Officer (Matthew Eby)

Road(s) and routes per MDRS Map: Hab area, mirror coating experiment sample location (close to Marble Ritual but not crossing Cow Dung Rd)

Mode of Travel: Walking

Commander Report – December 4th

Sol: 7
Summary Title: EVA #9 and Suit Troubleshooting
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 experienced another day of beautiful weather on Mars with mostly mild temperatures, sunny skies, and light high-altitude clouds. The crew enjoyed an egg, cheese, and vegetable casserole for breakfast. After a morning planning meeting, HSO collected crew health stats. CDR, Technology Officer, and Green Hab Officer conducted a short EVA in the early afternoon to deploy the ham radio antenna and fly two drone flights around the Hab and around the mirror coating experiment sample location. Fortunately, winds were light today, so the drone was able to successfully complete both flights. During the EVA, Engineer and HSO completed troubleshooting of the two-piece EVA suits and determined that the power strip the charging cables were plugged into had gone bad. They completed further tests with various combinations of suits, batteries, and chargers to ensure all six suits are now charging and holding charge properly. The crew relaxed a bit in the afternoon, reading, knitting, working on art projects, and sharing photos and stories. For dinner, the crew enjoyed tacos with homemade tortillas and baked a cake to celebrate the midpoint of their mission.

Operations Report – December 4th

SOL: 7

Name of person filing report: Ashley Kowalski

Non-nominal systems: N/A

Notes on non-nominal systems: N/A


Spirit rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: ~270 gallons remaining
Water (static tank): The crew was informed that one load of water resupply was delivered today (120 gallons). Yesterday, we were told we had 175 gallons left. Today, we estimate we used ~25 gallons. So, before the water resupply, we likely had ~150 gallons. Thus, once the resupply was delivered, we now have about ~270 gallons.
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The Crew Engineer and Health and Safety Officer performed the suggestions provided to the crew by Mission Support. We realized that the problem actually was with the power strip that all of the chargers were plugged into. This is the process of elimination that we went through in order to get to that conclusion:

We first opened EVA suits #3, #4, and #6 and inspected the wiring and the fuses. All of the wiring looked good. The Crew Engineer replaced the blown fuse in EVA Suit #4. We also confirmed that all of the chargers worked. They all read >13 V.

Next, we swapped batteries between suits. We took a battery from Suit #1, one of the known functional suits, and put it into each of the non-functioning suits (#3, #4, and #6). When we did this, each one of the suits turned on and the air flowed normally. Thus, we could conclude that there was nothing wrong with the suits themselves, but rather the problem lay with the batteries we took out of those suits.

Then, we put each of the batteries from the non-functioning suits into Suit #1 (the functioning suit) to see if those batteries would increase from their low values to a value >12 V. The interesting part here is that the battery would not increase its charge even in the good suit. In fact, it was losing charge. So, we still thought the problem must be with the battery.

However, we decided to try one last thing. We thought that perhaps the problem was with the power strip that all the chargers were plugged into. The lights on those chargers did not consistently shine red or green (depending on if the battery was charger or fully charged), so we plugged in EVA suits #3, #4, and #6 with their original batteries (the ones that were reading low voltage values) into a different power strip that was found on the table shelf directly under where the two-piece suits stand and left them to charge for a bit. When we came back, the battery for suit #3 read 13.3 V, the battery for suit #4 read 13.76 V, and the battery for suit #6 read 12.33 V.

Thus, we can confidently conclude that the power strip was the actual issue. All two-piece suits now have charged, are reading the correct voltage values, and are ready for use. Please let us know what you would like us to do with the power strip having issues. We have labeled it indicating that it is having issues.

Finally, the two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Additional information can be found in the GreenHab Report.
WATER USE: ~34 gallons (~3 gallons today)
Heater: On
Supplemental light: Lights were checked at 11:15 AM, and appear OFF as expected. The GreenHab Officer will check that they are on after 10 PM tonight.
Harvest: N/A

Summary of ScienceDome operations: No operations performed today by the crew in the Science Dome.
Dual split: N/A

Summary of RAM operations: The multimeter from the RAM was used to check the voltages of the two-piece EVA suits during today’s troubleshooting/inspections.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: The odor from the toilet area has dissipated. There is no longer an issue with smell.

Journalist Report – December 4th

Sol: 7
Author: Alli Taylor, Executive Officer
Title: Soup & Suits

After a Saturday night of trivia and belly laughs, the crew was able to sleep in for a more relaxed Sunday. I started the day by baking an experimental quiche casserole created with rehydrated ingredients including potato, egg, spinach, and Colby cheese. With ketchup, I found the dish quite edible and was surprised to find most of the crew seemed to enjoy it as well. Shortly after brunch we began suiting up Commander Kristine Ferrone, Technology Officer Trevor Jahn, and GreenHab Officer Matthew Eby for a short EVA around the vicinity of the Habitat. With the winds calm, Vixl the drone could be heard whirring through two pre-planned flight paths. While I ran IVA comms for the EVA crew, HSO Barbara Braun and Crew Engineer Ashley Kowalski could be heard troubleshooting spacesuits, carefully narrowing down possible root causes for anomalous battery charging to discover a faulty power strip. I was able to squeeze in a 15 minute cardio blast between comm check intervals as the trio outside the habitat flew Vixl on the pre-planned flight paths collecting a plethora of imagery. With a successful deployment of the ham radio antenna and completed drone flights, the crew returned from EVA. Ashley began cooking a late lunch of cheddar and broccoli soup, and the crew came together to share stories from experiences in EMT training and volunteer rescue. In the late afternoon the crew split up to work on individual reports, exercises, ham radio operations, and dinner prep. Meal time has definitely been something to look forward to as we work through our daily operations, with unjustifiably delicious outcomes thus far. The crew remains optimistic and generally in good spirits with regard to our mission, and I look forward to accomplishing our remaining goals and accumulating more lessons learned over the coming week.

GreenHab Report – December 4th

Crew 269 GreenHab Report 12-4-2022
GreenHab Officer: Matthew Eby
Environmental control: Heater on, door closed, 20% humidity.
Average temperatures: 65-deg F at time of watering. Range prior 24 hours: 65-deg F – 94-deg F
Hours of supplemental light: 4 hours
Daily water usage for crops: Approximately 5 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank: Approximately 210 gallons (Based on a 250 gallon capacity, visually 80-85% full)
Time(s) of watering for crops: 4:30 PM
Changes to crops: None
Narrative: Today continued to monitor the distressed pea plants, noted that soil has dried out after reducing watering in recent days, so a soaking was applied. Newer growth at the top of the plant looks healthy, see attached photo.
Harvest: None
Support/supplies needed: None needed at this time.

Copyright © The Mars Society. All rights reserved. | Main Site